Earlier today, we did a bunch of stuff to see if we could improve the situation with the excess ALS-X noise. Long story short, here are the parameters that were changed, and their initial and final values:
X-end laser diode temperature: 28.5 degrees ---> 31.3 degrees
X-end laser diode current: 1.900 A ---> 1.942 A
X-end laser crystal temperature: 47.43 degrees ---> 42.6 degrees
PSL crystal temperature: 33.43 degrees ---> 29.41 degrees
PSL Diode A temperature: 21.52 degrees ---> 20.75 degrees
PSL Diode B temperature: 22.04 degrees ---> 21.3 degrees
The Y-end laser temperature has not yet been adjusted - this will have to be done to find the Y-beatnote.
Unfortunately, this does not seem to have fixed the problem - I was able to find the beatnote, with amplitude on the network analyzer in the control room consistent with what we've been seeing over the last few days, but as is clear from Attachment 1, the problem persists...
Some details not directly related to this work: